5. Determine the z-score value in each of the following scenarios:
a. What z-score value separates the top 8% of a normal distribution from the bottom
92%?
b. What z-score value separates the top 72% of a normal distribution from the bottom
28%?
c. What z-score value form the boundaries for the middle 58% of a normal
distribution?
d. What z-score value separates the middle 45% from the rest of the distribution?
2. Annual salaries for a large company are approximately normally distributed with a mean of
R50,000 and a standard deviation of R20,000.
a. What salary would an employee need to get in order to be in the lowest 30%?
b. What is the probability of having an above average salary range of between R60000 to
R80000.
1
Expert's answer
2019-05-14T10:25:22-0400

a.

"P(Z<z^*)=0.92""z^*=1.4051"

b.


"P(Z<z^*)=0.28"

"z^*=-0.5828"

c.


"P(z^*_1<Z<z^*_2)=0.58"

"0.5-0.58\/2=0.21"

"P(Z<z^*_1)=0.21"

"z^*_1=-0.8064"

"z^*=\pm 0.8064"


So the interval is "(-0.8064, 0.8064)"  


d.


"0.5-0.45\/2=0.275"

"P(Z<z^*_1)=0.275"

"z^*_1=-0.5978"

"z^*=\pm 0.5978"

So the interval is "(-0.5978, 0.5978)."


2.


"\mu=R50,000""\sigma=R20,000"

a. What salary would an employee need to get in order to be in the lowest 30%? 

"Z={X-\mu \over \sigma}={X-\mu \over \sigma}"

"P(Z<z^*)=0.3=>z^*=-0.5244"

"x=z^*\sigma+\mu"

"x=-0.5244(R20000)+R50000=R39512"

b. What is the probability of having an above average salary range of between R60000 to

R80000.


"z_1={R60000-R50000 \over R20000}=0.5"

"z_2={R80000-R50000 \over R20000}=1.5"

"P(R60000<X<R80000)=P(0.5<Z<1.5)="

"=P(Z<1.5)-P(Z<0.5)="

"=0.93319280-0.69146246=0.24173034"